Please provide step by step instructions and print so I can read the response.

1) Find the threshold energy for the following reaction:

2) Calculate the energy released in the fission reaction:
Answer #1

1)

2 6 0 8

Mass of Oxygen Mo 15.99491 amu

Mass of neutron M, 1,008664 armu

Mass of carbon Mc 13.00335 amu

Mass of Helium MHe4.00260 amu

Threshold energy Mproducts+Mreactants ZİVItarget

R-(Mproducts - Mreactants)c

MproductsーMreactants)(Alproducts+ MTeactants 41Mtarget Mreactants Mgrofucts tMreactant 931 MeV th

931 Me th 2Mo

Eth 2.352 MeV

--------------------

Mass of Helium Мне 3.016029 amu

Mass of neutron M, 1,008664 armu

Mass of Hydrogen Мн = 2.0141020772 u

MproductsーMreactants)(Alproducts+ MTeactants 41Mtarget Mreactants Mgrofucts tMreactant 931 MeV th

Hel 931 MeV th 2MHe

E_{th}=4.54,MeV

---------------------

2)

Energy released in fission process 136 54 235 38

Mass of Uranium M_{U}=235.04393,amu

Mass of neutron M, 1,008664 armu

Mass of Strontium Msr87.905612 amu

Mass of Xenon Mxe 135.9072 19 a77221

Mass of products M_{products}=M_{Sr}+M_{Xe}+12M_n=235.9168,amu

Mass of reactants M_{reactants}=M_U+M_{n}+12M_n=236.0526,amu

Energy released is E=(M_{reactants}-M_{products})c^2=(236.0526-235.9168)*931=126.4,MeV